OpenStudy (anonymous):

Which number is a solution of the inequality? m > 13/3

OpenStudy (anonymous):

I assume this is multiple choice?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

A.3 B.5 C.2 D.-9

OpenStudy (anonymous):

Well the solution is going to be a number greater than 13/3. To determine if one of the choices is greater, you can rewrite the numbers in terms of thirds:\[3 \rightarrow\frac{9}{3}\]\[5 \rightarrow \frac{15}{3}\]\[2 \rightarrow \frac{6}{3}\]\[-9 \rightarrow -\frac{27}{3}\] So which one of those is greater than 13/3?